# A Non-perturbative General Approximation Scheme (NGAS) for Interacting Quantum Systems with Perturbation Theory for Arbitrary Strength of Interaction

### v2

@@ IX.3 The Effective Potential
[…]
@@
$<~{}{\cal H}_{0}~{}>=I_{1}(M)~{}-~{}3gI_{0}^{2}(M)~{}+~{}\frac{1}{2}~{}m^{2}\sigma^{2}~{}+~{}g~{}\sigma^{4}~{}~{}~{}$ (111)
@@
where, it is implicitly understood that the “gap-equation”, eqn.(101-102) is to be first solved to obtain ${}M^{2}~{}$as a function of ‘$\sigma$.
@@ IX.4 Equivalence to the Gaussian-Effective Potential
In view of the results derived in the previous subsection , the following remarks/ observations are in order:
(a)  The results contained in eqs. (101-102 ) for the gap equation and the integrals defined by eqs. (108) were first derived by Stevenson [18] in the context of the“Gaussian effective potential (GEP)” for the symmetric ${}g\phi^{4}~{}$ theory and obtained by variational calculation using a Gaussian-trial wave-function. The reproduction of the results of the GEP in ref.[18] in the LO of NGAS demonstrates that the GEP is contained in the NGAS as the leading order approximation.
(b)  In view of the established equivalence with the GEP, all the results obtained in the former approximation, are reproduced in the LO of the NGAS. In particular, the demonstration of non-triviality of the symmetric ${}g\phi^{4}~{}$ theory in the GEP,  being entirely based upon the consequences of renormalization is also reproduced in the LO of NGAS. We elaborate in the following sub-section, some of the properties of the theory concerning the stability and non-triviality of the theory.
(c) It must be emphasized, however, that the current scheme, NGAS is based upon entirely  different  starting  assumptions   and  is much more  general than the GEP, which is obtained solely due to the choice of the AH as given in eqn.(4.3) and that too, in the leading order.
@@ IX.5 Renormalization and realization of the ’non-triviality’ of the theory in the leading order of NGAS
Similarly,  after a straight forward calculation, one obtains ref.[18] the renormalized coupling as given by:
@@
An important consequence of eqn.(114) is that $g$ must be chosen negative: $g<0$, in order that the theory may be renormalized. This is because $I_{0}(m_{R})$ is divergent and the (unmeasurable) bare mass `$m^{\prime}$ of the theory may be infinite but the physical (renormalized) mass $m_{R}$ has to be finite. This version of the theory has therefore, been designated in ref.[18] as precarious ${}g\phi^{4}$ theory.
[…]
@@ IX.6 Emergence of the “Non-Triviality” of the theory
[…]
@@
[…]
@@
[…]
@@
where ${}\sigma^{2}~{}=~{}0~{}$ is again the location of the global-minimum of ${}U^{P}(\sigma)~{}$, as can be readily verified. In this context, it must be emphasized that the integrals : $\bar{I}_{n}$ occurring in eqn.(124) are $independent$ of $\sigma$.
@@
[…]
@@ IX.7 The Underlying Bogoliubov-Valatin Transformation
The actual/physical vacuum state in presence of interaction is approximated in the LO of NGAS by the state: $|~{}vac>$ which is the lowest energy state of ${}H_{0}~{}$. The structure and properties of this state can be inferred from studying the quantum-canonical transformation (i.e. Bogoliubov-Valatin transformation, ref.[46]) connecting the interacting vacuum state (IVS) with the free field vacuum ( FFV ) state.
[…]
@@
$[a(\vec{k}),~{}a^{\dagger}(\vec{q})]~{}=~{}\Omega_{k}(m)~{}\delta^{3}(\vec{k}-\vec{q}).$ (137)
@@
$[B(\vec{k}),B^{\dagger}(\vec{q})]~{}=~{}\delta^{3}(\vec{k}-\vec{q})~{}=~{}[A(\vec{k}),~{}A^{\dagger}(\vec{q}))].$ (140)
@@
To show this, consider eqs.(90) and (132) at ${}t~{}=~{}0~{}$, which can be written as:
@@ IX.8 Structure and Stability of the Effective Vacuum
where ${}\alpha^{vac}(\vec{k})~{}$ is given by eqn.(142) evaluated for ${}M~{}\equiv~{}m_{R}~{}$=  free-particle-mass renormalized  about the IVS, ${}|vac>$  (It may be recalled that ${}M(\sigma~{}=~{}0)~{}=~{}m_{R}~{}$ and ${}\sigma~{}=~{}0~{}$ define the IVS ). This leads finally to the expression :
@@ IX.9 Section-IX : Summary
We have demonstrated that the physical properties of the $\phi^{4}$ field theory depend critically on the structure of the physical vacuum state. The interacting vacuum state leads to a non-trivial and stable theory with finite renormalized parameters whereas, the physics developed around the non-interacting (‘perturbative’) vacuum becomes either unstable or trivial.
X Perturbation theory for arbitrary coupling Strength in NGAS - motivation and scope
We have discussed the key features of the standard formulation of perturbation theory (SFPT) in Section-I while surveying different approximation methods and noted its merits and the limitations. In this Section, we propose an improved perturbation theory, herein after designated as “Mean Field PerturbationTheory”(MFPT) which is based on NGAS and aimed at overcoming the limitations of SFPT. (The nomenclature will be explained in the sections to follow.) Some clues of how to achieve this objective have been provided at the end of Section-VIII and these will be further developed in this Section. In particular, we will explore how to preserve in MFPT, the nonlinearity and the analytic properties of the system-Hamiltonian and other “non-perturbative” effects , which are not achievable in SFPT.
However, there are certain features, tools and techniques used in SFPT which may be needed also in MFPT. In the following Section we consider those particular ’tools’ and ’techniques’ which may be relevant for this purpose.
@@ XI General features of the ‘standard’ formulation of perturbation theory(SFPT)
[…]
@@
where, the first term $\mathcal{E}_{0}$ represents the contribution known from the non-interacting theory while $\mathcal{E}_{k}$ denotes the perturbative correction (PC) at the $k$-th order. In order that the PC here make sense it is ab initio necessary that the coupling strength, $g$ be sufficiently small, i.e. ${}g\ll 1$.
[…]
@@
where $a,b,c$ are constants depending upon the particular theory under consideration. Moreover, the asymptotic-nature of the PS manifests through the following property [107,108]:
@@
In that context, historically the systematic investigations of large-order behavior of (renormalized) perturbation expansion were primarily established by studies [88] of systems with anharmonic interactions (AHI) in quantum mechanics (QM) and through subsequent investigations [107] in other systems.
It is now established, (see, e.g.[98,107,108]) that generalized-factorial growth of perturbation correction is a general feature of perturbation theory for almost all cases of physical interest and further that this behavior is primarily due to the factorial growth of Feynman diagrams/ intermediate states contributing at large orders irrespective of the theory under consideration.
Therefore, It becomes foremost necessary, to assign a meaning to the asymptotic-perturbation theory and investigate whether the latter can be “summed” in some sense to yield a finite result. We will focus on two main methods to achieve this goal, namely the method of “Optimal Truncation of the Asymptotic Series” and the “method of Borel-summation”.These are discussed in some detail in the next subsection.
@@ XI.1 Making sense of divergent, asymptotic series: Optimal truncation
This method utilizes the property [107,108], that the initial terms of an AS continue to decrease in magnitude till the term of the least magnitude(TLM) is reached, beyond which the subsequent terms exhibit monotonic increase. Hence, the AS can be reasonably approximated by truncation at the TLM.
XI.2 Making Sense of Divergent, Asymptotic Series: Borel-Resummation
@@
In view of the inadequacies of SFPT mentioned in Section-II and the previous section, it is perhaps imperative to explore whether perturbation theory could at all be liberated from the limitations of the small-coupling-power-series-expansion. The following Sections devoted to such an exploration in some detail.
@@ XII General Features of the ‘Mean Field Perturbation Theory’ (MFPT) in NGAS
One of the main motivations for proposing the NGAS as described earlier, is the possibility of construction of an improved perturbation theory which could be convergent/summable for all allowed values of the coupling strength, $g$.
Working in the standard Hamiltonian formalism, we can construct the improved perturbation theory in which the system-Hamiltonian, $H(g)$ can be split in to an exactly solvable dominant part $H_{0}$ and a sub-dominant perturbation $H^{\prime}$, i.e.
@@
Then, for the realization of a perturbative-framework for arbitrary values of the coupling $g$, it may be sufficient to fulfill the following two conditions:
(a) both $H_{0}$ and $H^{\prime}$ must depend non-trivially upon the coupling strength $g$, i.e. $H_{0}=H_{0}(g)$ and $H^{\prime}=H^{\prime}(g)$ such that the dominant $g$-dependence resides in $H_{0}(g)$, and further that (b) the contribution of the perturbing Hamiltonian $H^{\prime}(g)$ remains sub-dominant for arbitrary value of $g$.
As discussed earlier ( Section-VI), a practical implementation of the above conditions can be achieved [82,104] through the following steps:
(i) choose $H_{0}$ to depend upon a suitable set of free parameters ${{\alpha_{i}}}$: $H_{0}=H_{0}(\{\alpha_{i}\})$,
(ii) impose the constraint that
@@
where,$|\phi_{n}\rangle$ is defined by the eigen-value equation for $H_{0}$ given by
@@
with ``$n"$ denoting the spectral-label,
(iii)finally, determine the parameters $\{\alpha_{i}\}$ by variational minimization of $\langle H(g)\rangle$ and subsequent imposition of the constraint given by eqn.(154).
By this simple procedure, the non-linear $g$-dependence of the system-Hamiltonian, $H(g)$ is fed back into the reference Hamiltonian:
@@
Solving the eigen-value equation (eqn.155) for $H_{0}(g)$, then leads to [82,104] the results for $E_{0}(g)$ which, constitute the leading-order (LO) approximation. These LO-results are found [82,104] to be quite accurate uniformly over all spectral-labels ‘$n$’and for arbitrary physical values of $g$. It may be noted that this procedure is non-perturbative, self-consistent and preserves the analytic-structure of $E(g)$ in the $g$-plane [82,104].
The above procedure has already been discussed under Section-VI but succinctly recapitulated here for subsequent discussions.
Thus, having achieved the incorporation of the major $g$-dependence at the LO, the task remains to build a perturbation theory about the approximating-Hamiltonian (AH), $H_{0}$ and systematically improve the accuracy still further by computing order-by-order, the residual corrections as described in the next subsection.

### v3

@@ IX.3 The Effective Potential
It may be noted that the l.h.s. of eqn. (103) is $\it{defined}~$ to be a function of $\sigma~$ alone. This means that any other parameter occuring in ${\cal H}_{0}~$ is to be fixed by variational-minimization of ${\cal H}_{0}>$. The procedure is made explicit below by working out the current example, with ${\cal H}_{0}~$ defined in eqn.(86). To this end, we first calculate ${\cal H}_{0}~>~$ which guarantees the following equation:
@@
${\cal H}_{0}~>=I_{1}(M)~-~3gI_{0}^{2}(M)~+~\frac{1}{2}~m^{2}\sigma^{2}~+~g~\sigma^{4}~~~$ (111)
@@
where, it is implicitly understood that the “gap-equation”, eqn.(101-102) is to be first solved to obtain $M^{2}~$as a function of $\sigma$.
@@ IX.4 Equivalence to the Gaussian-Effective Potential
In view of the results derived in the previous subsection , the following remarks/ observations are in order: (a)  The results contained in eqs. (101-102 ) for the gap equation and the integrals defined by eqs. (108) were first derived by Stevenson [18] in the context of the“Gaussian effective potential (GEP)” for the symmetric $g\phi^{4}~$ theory and obtained by variational calculation using a Gaussian-trial wave-function. The reproduction of the results of the GEP in ref.[18] in the LO of NGAS demonstrates that  the GEP is contained in the NGAS as the leading order approximation. (b)  In view of the established equivalence  with the GEP, all the results obtained in the former approximation, are reproduced in the LO of the NGAS. In particular, the demonstration of non-triviality of the symmetric $g\phi^{4}~$ theory in the GEP,  being entirely based  upon the consequences of renormalization  is also reproduced in the LO of NGAS. (c) It must be emphasized, however, that the current scheme, NGAS is based upon entirely  different  starting  assumptions   and  is much more  general than the GEP, which is obtained solely due to the choice of the AH as given in eqn.(4.3) and that too, in the leading order.
We elaborate in the following sub-section, some of the properties of the theory  concerning the stability  and non-triviality of the theory.
@@ IX.5 Renormalization and realization of the ’non-triviality’ of the theory in the leading order of NGAS
Similarly,  after a straight forward calculation, one obtains [18] the renormalized coupling as given by:
@@
An important consequence of eqn.(114) is that $g$ must be chosen negative: $g<0$, in order that the theory may be renormalized. This is because   $I_{0}(m_{R})$ is divergent and the (unmeasurable) bare mass $m^{\prime}$ of the theory may be infinite but the physical (renormalized) mass $m_{R}$ has to be finite. This version of the theory has therefore, been designated in ref.[18] as    precarious $g\phi^{4}$   theory.
We next discuss some of the consequences of the above non-perturbative renormalization scheme obtained in the LO of NGAS leading to the $\it{stability}$ and $\it{non-triviality}$ of the theory.
@@ IX.6 Emergence of the “Non-Triviality” of the theory
$+\mbox{(sub-leading~ terms)}$ (119)
With the aid of eqs.(116) and (117),  the effective potential, as given by eqn.(110) can be recast in manifestly renormalized form involving the $\it{observable}~$ parameters: $g_{R}~$ and $m_{R}~$ only. The resulting expression is given by :
@@
It must be pointed out that, one has to $\it{first}~$ solve the gap-equation, eqn.(121) to obtain $t~\equiv~t(\sigma)~$, which is then to be substituted in eqn.(118) to infer the $\sigma-$dependence of $U_{0}(\sigma)~$.
@@
The domain of validity for the effective potential (EP) is constrained by the range of $\sigma^{2}~$ as determined by equation (122) for any given value of $\eta$. Specifically, in the regime of large coupling ($\eta~\rightarrow~0~$), the domain of the EP diminishes with decreasing $\eta~$, as $\sigma^{2}{min}~\rightarrow 0$ in this limit. This scenario corresponds to small oscillations around $\sigma~\simeq~0~$, which is indicative of a pathological situation when $|g{R}|~\rightarrow$$\infty$. Conversely, in the small coupling regime ($g_{R}~\rightarrow~0$), which corresponds to $|\eta|>>~1$, the domain of the EP expands with increasing $\eta~$. Thus, it can be summarized that the leading-order effective potential (LO-EP) of the symmetric $g\phi^{4}~$ theory is reasonable and well-behaved unless the renormalized coupling is excessively large. Consequently, one can conclude that a non-trivial and stable theory emerges in the leading order of NGAS provided the $\it{physical}~$ coupling $g_{R}~$ is $\it{not}~$ unusually large.
Further commentary on the stability of the perturbative vacuum of the theory is as follows. To examine the stability issue, it is necessary to compute the effective potential based on the $\it{perturbative}~$ vacuum (i.e., the vacuum of the free-field theory). This is achieved by setting $M~\rightarrow~m~$. Thus, starting from equation (106) and letting $M~\rightarrow~m~$, one obtains, after simplification, the following expression:
@@
where $\sigma^{2}~=~0~$ is again the location of the global-minimum of $U^{P}(\sigma)~$, as can be readily verified. In this context, it must be emphasized that the integrals : $\bar{I}_{n}$ occurring in eqn.(124) are independentindependent of $\sigma$.
@@
The requirement for the finiteness of $\bar{m}_{R}~$ and $\bar{g}_{R}~$ necessitates that $g$ must be negative (refer to equation (127)), for otherwise $\bar{m}_{R}^{2}~$ would be infinitely large since $\bar{I}_{0}~$ is divergent and the bare (unobservable) mass, $m^{2}>0$. However, this would lead to instability since the effective potential $U^{P}(\sigma)~$ will not have a lower bound. This is made manifest by explicitly writing the EP in terms of the renormalized parameters:
@@ IX.7 The Underlying Bogoliubov-Valatin Transformation
The actual/physical vacuum state in presence of interaction is approximated in the LO of NGAS by the state: |vac>|~vac> which is the lowest energy state of $H_{0}~$. The structure and properties of this state can be inferred from studying the quantum-canonical transformation (i.e. Bogoliubov-Valatin transformation, ref.[46]) connecting the interacting vacuum state (IVS) with the free field vacuum ( FFV ) state.
For this purpose it is convenient to start from the Fourier-decomposition of the field $\phi(\vec{x},t)~$ in terms of the $\it{free~field}~$ creation-and annihilation operators analogous to eqn.(88):
@@
$a(\vec{k}),~a^{\dagger}(\vec{q})]~=~\Omega_{k}(m)~\delta^{3}(\vec{k}-\vec{q}).$ (137)
@@
$B(\vec{k}),B^{\dagger}(\vec{q})]~=~\delta^{3}(\vec{k}-\vec{q})~=~[A(\vec{k}),~A^{\dagger}(\vec{q}))].$ (140)
@@
To show this, consider eqs.(90) and (132) at t=0~t~=~0~, which can be written as:
@@ IX.8 Structure and Stability of the Effective Vacuum
where $\alpha^{vac}(\vec{k})~$ is given by eqn.(142) evaluated for $M~\equiv~m_{R}~$=  free-particle-mass renormalized  about the IVS, |vac>~|vac>  (It may be recalled that $M(\sigma~=~0)~=~m_{R}~$ and $\sigma~=~0~$ define the IVS ). This leads finally to the expression :
@@ IX.9 Section-IX : Summary
We have demonstrated that application of NGAS to the ${\lambda}{\phi}^{4}$ quantum field theory under the harmonic approximation, reproduces the main results of the Gaussian-effective potential (GEP) method . In that sense, the scope of NGAS is broader since the GEP is contained as a particular realization of the latter. Beyond that, NGAS reveals the condensate structure of the effective vacuum and highlights the instability of the perturbative ground state. Besides, the effective vacuum state in NGAS leads to a non-trivial and stable theory with finite renormalized parameters whereas, the physics developed around the non-interacting (‘perturbative’) vacuum becomes either unstable or trivial.
X Perturbation Theory for Arbitrary Coupling Strength in NGAS: Motivation and Scope
In Section-I, we have examined the principal characteristics of the standard formulation of perturbation theory (SFPT) while reviewing various approximation methods, noting both its advantages and limitations. In this section, we introduce an enhanced perturbation theory, hereafter referred to as ”Mean Field Perturbation Theory” (MFPT), which is based on NGAS and aims to address the shortcomings of SFPT. (The terminology will be elucidated in subsequent sections.) Preliminary insights into achieving this objective were provided at the conclusion of Section-VIII, and these will be further developed in this Section.
Specifically, we will investigate how MFPT can maintain the nonlinearity and analytic properties of the system Hamiltonian, as well as other ”non-perturbative” effects, which are not attainable in SFPT. However, certain features, tools, and techniques employed in SFPT may also be necessary for MFPT. In the following Section, we will consider those specific ’tools’ and ’techniques’ that may be pertinent for this purpose.
@@ XI General features of the ‘standard’ formulation of perturbation theory(SFPT)
As highlighted in Section-II, perturbation theory remains the preferred method of approximation for various practical and theoretical reasons [9]. However, SFPT is subject to several limitations, primarily due to its $\it defining$} characteristic that results for physical observables are expressed as a power-series expansion in the coupling strength (CS). To illustrate, consider the energy $E(g)$ as the observable, represented by the standard expansion given by eqn.(1), i.e.,
@@
where the first term $\mathcal{E}_{0}$ represents the contribution known from the non-interacting theory, while $\mathcal{E}_{k}$ denotes the perturbative correction (PC) at the $k$-th order. For the PC to be meaningful, it is ab initio necessary that the coupling strength, $g$, be sufficiently small, i.e., $g\ll 1$. Nevertheless, even when $g~$ is appropriately small, it is generally observed [107] that SFPT results in $\it divergent$} but $\it asymptotic$} perturbation series (PS). Specifically, the PS exhibits [107] (generalized) factorial growth at large order, as described by the following generic form:
@@
where $a,b,c$ are constants depending on the particular theory under consideration. Moreover, the asymptotic nature of the PS is manifested through the following property [107,108]:
@@
In this context, the systematic investigations of the large-order behavior of (renormalized) perturbation expansion were historically established through studies [88] of systems with anharmonic interactions (AHI) in quantum mechanics (QM) and subsequent investigations [107] in other systems. It is now established (see, e.g., [98,107,108]) that the generalized-factorial growth of perturbation correction is a general feature of perturbation theory for almost all cases of physical interest, and further, that this behavior is primarily due to the factorial growth of Feynman diagrams/intermediate states contributing at large orders, irrespective of the theory under consideration. Therefore,  It becomes foremost necessary, to   assign a meaning to the asymptotic-perturbation theory and investigate whether the latter can be “summed” in some sense to yield a finite result. We will focus on two main methods to achieve this goal, namely the method of “Optimal Truncation of the Asymptotic Series” and the “method of Borel-summation”.These are discussed in some detail in the next subsection.
@@ XI.1 Making sense of divergent, asymptotic series: Optimal truncation
This method utilizes the property [107,108], that the initial terms of an AS continue to decrease in magnitude till the term of the least magnitude(TLM) is reached, beyond which the subsequent terms exhibit monotonic increase ( refer to the defining equation, eqn.(152)). Hence, the AS can be reasonably approximated by truncation at the TLM.
XI.2 Making Sense of Divergent, Asymptotic Series: Borel-Summation
@@
In light of the deficiencies of SFPT discussed in Section-II and the preceding section, it is essential to investigate whether perturbation theory can be freed from the constraints of the small-coupling-power-series-expansion. The subsequent Sections are dedicated to this exploration in detail.
@@ XII General Features of the ‘Mean Field Perturbation Theory’ (MFPT) in NGAS
One of the primary motivations for proposing the NGAS, as previously described, is the potential to develop an enhanced perturbation theory that is convergent or summable for all permissible values of the coupling strength, $g$.
Utilizing the standard Hamiltonian formalism, we can construct an improved perturbation theory wherein the system-Hamiltonian, $H(g)$, is divided into an exactly solvable dominant component $H_{0}$ and a sub-dominant perturbation $H^{\prime}$, such that
@@
To establish a perturbative framework for arbitrary values of the coupling $g$, it may be sufficient to satisfy the following two conditions: (a) both $H_{0}$ and $H^{\prime}$ must depend non-trivially on the coupling strength $g$, i.e., $H_{0}=H_{0}(g)$ and $H^{\prime}=H^{\prime}(g)$, with the dominant $g$-dependence residing in $H_{0}(g)$, and (b) the contribution of the perturbing Hamiltonian $H^{\prime}(g)$ remains sub-dominant for any value of $g$.
As previously discussed (Section-VI), a practical implementation of these conditions can be achieved [82,104] through the following steps: (i) select $H_{0}$ to depend on a suitable set of free parameters $\alpha_{i}$}}: $H_{0}=H_{0}(\{\alpha_{i}\})$, (ii) impose the constraint that
@@
where $|\phi_{n}\rangle$ is defined by the eigenvalue equation for $H_{0}$ given by
@@
with n``n" denoting the spectral label, (iii) finally, determine the parameters $\{\alpha_{i}\}$ by variational minimization of $\langle H(g)\rangle$ and subsequent imposition of the constraint given by eqn.(154).
Through this straightforward procedure, the non-linear $g$-dependence of the system-Hamiltonian, $H(g)$, is incorporated into the reference Hamiltonian:
@@
Solving the eigenvalue equation (eqn.155) for $H_{0}(g)$ then yields [82,104] the results for $E_{0}(g)$, which constitute the leading-order (LO) approximation. These LO results are found [82,104] to be highly accurate across all spectral labels ‘$n$’ and for any physical values of $g$. It is noteworthy that this procedure is non-perturbative, self-consistent, and preserves the analytic structure of $E(g)$ in the $g$-plane [82,104].
The aforementioned procedure has been previously discussed in Section-VI but is briefly recapitulated here for subsequent discussions. Having incorporated the major $g$-dependence at the LO, the task remains to develop a perturbation theory around the approximating Hamiltonian (AH), $H_{0}$, and systematically enhance the accuracy by computing order-by-order residual corrections, as described in the next subsection.